# Self-reinforcing cascades: A spreading model for beliefs or products of varying intensity or quality

### v1

@@ Discussion
The self-reinforcing cascade process presents key features that makes it particularly appealing for modeling contagions observed in socio-technical systems. It is a parsimonious model to capture the fact that the strength of individual beliefs or the quality of products may vary and influence the ability of an individual to further transmit the cascade. This variability is aligned with real-world phenomena, where not all individuals or contents are equally influential in the transmission of ideas or behaviors. With this simple mechanism, the SRC model can produce a wide range of scaling behaviors for cascade size distributions, whereas classic percolation is constrained by a unique and universal scaling exponent obtained only at a precise critical point.
While self-similar or isotropic spreading rules are mathematically convenient, providing straightforward solutions and modeling frameworks, self-reinforcing cascades can offer advantages that warrant further study. The flexibility and richness of their outcome suggest that they are better suited to capture the complexities of real-world social contagions.
We outlined several important properties of self-reinforcing cascades and proposed three analytical approaches—exact probability generating functions, explicit solution of expected cascades, and the traveling-wave technique—to better understand these processes. This model may provide a useful framework for researchers and practitioners seeking to understand cascading behavior in complex real-world systems.
@@ Acknowledgments
The authors acknowledge financial support from The National Science Foundation awards #2419733 (L.H.-D. and G.B.) and #2242829 (J.L. and G.B.) as well as from Science Foundation Ireland under Grant number 12/RC/2289 P2 (J.P.G).

### v2

@@ Discussion
The self-reinforcing cascade process is a parsimonious model to capture the fact that the strength of individual beliefs or the quality of products may vary and thus influence their ability to spread further. This variability is aligned with real-world phenomena, where not all individuals or contents are equally influential in the transmission of ideas or behaviors. With this simple mechanism, the SRC model can produce a wide range of scaling behaviors for cascade size distributions, whereas classic cascade models are constrained by a unique and universal scaling exponent obtained only at a precise critical point.
The shape and statistics of SRC are surprising in two ways. One, they follow fat-tailed cascade size distribution in their subcritical regime. Two, the tail of their depth distribution remains exponential even at the critical point. Thus, subcritical and critical SRC are short but very broad compared to classic cascade models. The unique shape of SRC could help identify signatures of varying cascade intensity or quality in empirical data. We could expect this mechanism to appear in cascades where content is able to change (e.g., mutations in epidemics or personalized social media posts) as opposed to cascade with fixed content (e.g., cascade of clones, or reposting in social media), and SRC could provide more realistic models in these cases. In addition, the search remains open for a general model that provides anomalous exponents for both size and depth distributions as found empirically Notarmuzi et al. (2022).
We outlined several important properties of self-reinforcing cascades and proposed a few analytical approaches to better understand these processes: exact probability generating functions for cascade size, exact recursions over cascade depth, and the traveling wave technique for cascade intensity. Combining all three approaches, we were able to characterize the different scaling behaviors produced by the model. Altogether, this effort may provide a useful framework for researchers and practitioners seeking to understand cascading behavior in complex real-world systems.
@@ Acknowledgments
The authors acknowledge financial support from The National Science Foundation awards #2419733 (L.H.-D. and G.B.) and #2242829 (J.L. and G.B.) as well as from Science Foundation Ireland under Grant number 12/RC/2289 P2 (J.P.G.). The research was supported by International Partnerships and Programs at the University of Vermont as a Global Catalyst Research Partnership Award.